OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -6 and a vertex at (-8, -8) Format using this sample answer: f(x)=(3)(x^2-21x-19)

OpenStudy (eyust707):

well the way i would do it would be: if it has roots -10 and -6 we know that when you factor it it looks like: (x+6)(x+10) which equals x^2 +16x+60 and if the vertex is at -8,-8 then -b/2a = -8 right now -b/2a is -16/2 which ='s -8 so thats good! then lets check the y coordinate of the vertex our equation is y = x^2+16x + 60 and if we plug in a "-8" we get: y = (64) - 128 + 60 = -4 darn well thats half so lets times the whole thing by 2 2(x+6)(x+10)

OpenStudy (eyust707):

check that through the same process and sure enough it works!!

OpenStudy (eyust707):

i feel like there may be a simpler way... but I am not sure what it would be